OpenStudy (anonymous):

***MEDAL*** What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374?

OpenStudy (campbell_st):

well the sum of an arithmetic sequence is \[S_{n} = \frac{n}{2}[2a + (n - 1) \times d]\] you know a = -13 and n = 43 and the sum is 9374 just substitute them and solve for d.

OpenStudy (anonymous):

thanks!